# DeepProtein: Deep Learning Library and Benchmark for Protein Sequence Learning

### v1

@@ 3.3 Experimental Setup and Implementation Details
Code Base.
This library is an extension of the well-established drug discovery library, DeepPurpose (Huang et al., 2020), building upon its foundational capabilities to offer enhanced features for protein-related tasks. By leveraging the strengths of DeepPurpose, this new library provides additional tools and functionalities tailored specifically for protein science. The library is publicly available at https://github.com/jiaqingxie/DeepProtein/tree/main.
Hardware Configuration.
All experiments that are mentioned in this paper were trained on a 40GB NVIDIA A40 and a 24GB NVIDIA RTX 3090. The parameters we provide have ensured the scalable training on these two types of GPUs. When running GNNs on protein localization tasks, we observed a large portion of GPU memory occupied irregularly, so we recommend cutting down the size of the number of workers from 8 to 4 or batch size from 32 to 8 or even smaller to potentially avoid GPU out-of-memory (OOM) problems.
Software Configuration.
The library is implemented in Python 3.9, PyTorch 2.3.0, PyTDC 0.4.1 (Huang et al., 2021), DeepPurpose 0.1.5 (Huang et al., 2020), and RDKit 2023.9.6 (Landrum et al., 2006), scikit-learn 1.2.2 (Pedregosa et al., 2011), and DGLlife 0.3.2 (Li et al., 2021a). Besides, wandb is included in DeepProtein so that researchers can observe the visualization of training curves and test results easily.
Figure 2: Performance of selected models including CNN, Transformer, GCN and NeuralFP on SubCellular, PPI Affinity and CRISPR Leenay. We observe that CNN performs significantly better than GCN.
Table 2: Results of protein function prediction. The $\uparrow$ symbol indicates that higher values are better for the corresponding metric. For each method, we employed five different random seeds to perform independent runs, reporting the average results along with their standard deviations. On each task, the best method is bolded, and the second best is underlined. We use “**” to denote the method that achieves statistically better results than all the other methods (through statistical tests).
Model Fluorescence ($\rho$uparrow) Stability ($\rho$uparrow) $\beta$-lactamase ($\rho$uparrow) Solubility (PR-AUC $\uparrow$)
@@
CNN 0.680 $\pm$ 0.001 0.715 $\pm$ 0.025** 0.721 $\pm$ 0.020** $74.61\pm 0.55$
CNN-RNN $0.678\pm 0.001$ $0.638\pm 0.021$ $0.695\pm 0.012$ $74.10\pm 1.48$
Transformer $0.648\pm 0.001$ $0.442\pm 0.030$ $0.348\pm 0.002$ 78.86 $\pm$ 0.46**
GCN $0.397\pm 0.002$ $0.392\pm 0.027$ $0.417\pm 0.001$ $69.26\pm 0.73$
GAT $0.251\pm 0.002$ $0.165\pm 0.026$ $0.196\pm 0.013$ $62.44\pm 0.01$
NeuralFP $0.413\pm 0.011$ $0.333\pm 0.041$ $0.133\pm 0.020$ $78.74\pm 0.24$
AttentiveFP $0.260\pm 0.003$ $0.254\pm 0.022$ $0.058\pm 0.011$ $78.39\pm 0.19$
MPNN $0.237\pm 0.020$ $0.110\pm 0.021$ $0.068\pm 0.015$ $62.53\pm 0.31$
PAGTN $0.188\pm 0.036$ $0.266\pm 0.016$ $0.092\pm 0.018$ $61.33\pm 0.91$
Graphormer $0.067\pm 0.002$ OOM $0.103\pm 0.018$ OOM
@@
Model Subcellular (Acc $\uparrow$) Binary (PR-AUC $\uparrow$)
@@
CNN 50.18 $\pm$ 1.21 $90.88\pm 0.31$
CNN-RNN 52.58 $\pm$ 0.11** 91.43 $\pm$ 0.45**
Transformer $42.63\pm 0.68$ $78.38\pm 0.25$
GCN $47.45\pm 0.47$ $83.43\pm 0.10$
GAT $45.14\pm 0.10$ $82.15\pm 0.41$
NeuralFP $45.20\pm 0.49$ $81.14\pm 0.06$
AttentiveFP $42.38\pm 1.25$ $80.58\pm 0.30$
@@
Model PPI Affinity ($R^{2}$uparrow) Yeast (PR-AUC $\uparrow$) Human PPI (PR-AUC $\uparrow$)
@@
CNN $0.493\pm 0.015$ $51.93\pm 0.92$ $70.37\pm 1.22$
CNN-RNN 0.584 $\pm$ 0.026** $53.28\pm 0.85$ $70.45\pm 2.68$
Transformer $0.425\pm 0.021$ $53.79\pm 1.07$ $59.36\pm 4.00$
GCN $0.394\pm 0.006$ 58.98 $\pm$ 0.72 82.21 $\pm$ 1.13
GAT $0.230\pm 0.001$ $53.72\pm 0.39$ $77.63\pm 3.13$
NeuralFP $0.100\pm 0.054$ $57.00\pm 1.51$ $80.11\pm 1.25$
Table 5: Results of epitope and paratope prediction (residue-level classification).
Model IEDB (ROC-AUC $\uparrow$) PDB-Jespersen (ROC-AUC $\uparrow$) SAbDab-Liberis (ROC-AUC $\uparrow$)
@@
Model TAP (MAE $\downarrow$) SAbDab-Chen (MAE $\downarrow$) CRISPR-Leenay (MAE $\downarrow$)
@@
CNN $3.217\pm 0.026$ 0.219 $\pm$ 0.006 0.0745 $\pm$ 0.0005
CNN-RNN 0.712 $\pm$ 0.069** $0.226\pm 0.001$ $0.0755\pm 0.0010$
Transformer $3.476\pm 0.004$ $0.238\pm 0.012$ $0.1216\pm 0.0020$
GCN $2.761\pm 0.155$ $0.326\pm 0.015$ $0.1274\pm 0.0019$
GAT $2.675\pm 0.022$ $0.310\pm 0.010$ $0.1232\pm 0.0001$
NeuralFP $3.436\pm 0.015$ $0.253\pm 0.011$ $0.1243\pm 0.0001$
@@ 3.4 Results & Analysis
For each method, we used five different random seeds to conduct independent runs and reported the average results and their standard deviations. The results of protein function prediction are reported in Table 2.
@@
Key Observations.
We summarize the following key observations as takeaways.
  • Sequence-based neural architectures are powerful compared with graph-based neural architectures. Sequence-based neural architectures, such as CNN, RNN, and transformer, obtain superior performance in most protein sequence learning tasks. Specifically, in 12 out of all the 15 tasks across various protein sequence learning tasks, the sequence-based model (CNN, RNN, and transformer) takes the top-2 position.
  • Among all the 12 GNN-solvable tasks (except residue-level classification), graph neural networks (GNN) obtain the best accuracy among all the methods only in two protein-protein interaction (PPI) tasks, including Yeast and Human PPI.
  • Among all the graph neural networks (GNNs) across the whole 12 GNN-solvable tasks (except residue-level classification), the earliest variant, GCN (Kipf and Welling, 2016), achieves the best performance in 9 tasks.
  • Stability. From the learning curve (Figure 7-10), we find that GNN’s training curve is not stable. In contrast, the sequence-based models, including CNN, RNN, and transformer, converge more stably from the learning curve. This can be observed from Figure 7-9. On the contrary, training is more stable, fast and accurate for GAT when it comes to TAP dataset.
  • Computational complexity. The runtime and memory costs are reported in Figure 35 and 6. We find that GNN-based models are typically computationally inefficient. The key reason behind this is that GNN utilizes molecular graph as the feature, where each atom corresponds to a node and each chemical bond corresponds to an edge. While another model, such as CNN, RNN, and transformer, uses amino acid sequences as the input feature.
Figure 3: Empirical computational complexity on CRISPR and SAbDab Chen measured by runtime and memory.

### v2

@@ 3.3 Experimental Setup and Implementation Details
Code Base. This library is an extension of the well-established drug discovery library, DeepPurpose (Huang et al., 2020), building upon its foundational capabilities to offer enhanced features for protein-related tasks. By leveraging the strengths of DeepPurpose, this new library provides additional tools and functionalities tailored specifically for protein science. The library is publicly available at https://github.com/jiaqingxie/DeepProtein/.
Hardware Configuration. All experiments that are mentioned in this paper were trained on a 40GB NVIDIA A40 and a 24GB NVIDIA RTX 3090. For DeepProt-T5 fine-tuning, two 24GB NVIDIA RTX 3090 were used. The parameters we provide have ensured the scalable training on these two types of GPUs. When running GNNs on protein localization tasks, we observed a large portion of GPU memory occupied irregularly, so we recommend cutting down the size of the number of workers from 8 to 4 or batch size from 32 to 8 or even smaller to potentially avoid GPU out-of-memory (OOM) problems.
Software Configuration. The library is implemented in Python 3.9, PyTorch 2.3.0, PyTDC 0.4.1 (Huang et al., 2021), DeepPurpose 0.1.5 (Huang et al., 2020), and RDKit 2023.9.6 (Landrum et al., 2006), scikit-learn 1.2.2 (Pedregosa et al., 2011), and DGLlife 0.3.2 (Li et al., 2021a). Besides, wandb is included in DeepProtein so that researchers can observe the visualization of training curves and test results easily. More details about environment setup could be found in the GitHub.
Table 2: Results of protein function prediction. The $\uparrow$ symbol indicates that higher values are better for the corresponding metric. For each method, we employed five different random seeds to perform independent runs, reporting the average results along with their standard deviations. On each task, the best method is bolded, and the second (or the third closest to the second) best is underlined. We use “**” to denote the method that achieves statistically better results than all the other methods (through statistical tests).
Model Fluorescence Stability $\beta$-lactamase Solubility
$\rho$uparrow MAE $\downarrow$ $\rho$uparrow MAE $\downarrow$ $\rho$uparrow MAE $\downarrow$ PR-AUC $\uparrow$ Averaged F1 $\uparrow$
@@
CNN 0.680 $\pm$ 0.001 0.194 $\pm$ 0.001 0.715 $\pm$ 0.025 0.312 $\pm$ 0.008 0.721 $\pm$ 0.020 0.152 $\pm$ 0.001 $74.61\pm 0.55$ 62.50 $\pm$ 2.69
CNN-RNN $0.678\pm 0.001$ 0.268 $\pm$ 0.002 $0.635\pm 0.025$ 0.323 $\pm$ 0.001 $0.695\pm 0.012$ 0.157 $\pm$ 0.002 $75.46\pm 0.03$ 60.87 $\pm$ 3.19
Transformer $0.648\pm 0.001$ 0.371 $\pm$ 0.005 $0.375\pm 0.052$ 0.411 $\pm$ 0.026 $0.310\pm 0.041$ 0.232 $\pm$ 0.007 78.86 $\pm$ 0.46 61.28 $\pm$ 1.42
GCN $0.397\pm 0.002$ 0.864 $\pm$ 0.018 0.443 $\pm$ 0.017 0.445 $\pm$ 0.011 $0.450\pm 0.006$ 0.225 $\pm$ 0.001 $69.26\pm 0.73$ 56.30 $\pm$ 1.39
GAT $0.249\pm 0.001$ 1.201 $\pm$ 0.005 $0.101\pm 0.001$ 0.740 $\pm$ 0.002 $0.196\pm 0.013$ 0.264 $\pm$ 0.002 $62.44\pm 0.01$ 47.79 $\pm$ 0.42
NeuralFP $0.349\pm 0.049$ 1.016 $\pm$ 0.081 $0.373\pm 0.023$ 0.484 $\pm$ 0.020 $0.171\pm 0.007$ 0.263 $\pm$ 0.001 $78.74\pm 0.24$ 62.72 $\pm$ 1.91
AttentiveFP $0.180\pm 0.006$ 1.285 $\pm$ 0.001 $0.013\pm 0.004$ 0.763 $\pm$ 0.003 $0.058\pm 0.011$ 0.262 $\pm$ 0.007 60.56 $\pm$ 0.74 39.21 $\pm$ 0.66
MPNN $0.249\pm 0.002$ 1.275 $\pm$ 0.011 $0.118\pm 0.061$ 0.732 $\pm$ 0.006 $0.068\pm 0.015$ 0.262 $\pm$ 0.003 $62.53\pm 0.31$ 40.56 $\pm$ 5.46
PAGTN $0.139\pm 0.014$ 1.255 $\pm$ 0.040 $0.088\pm 0.194$ 0.727 $\pm$ 0.011 $0.092\pm 0.018$ 0.255 $\pm$ 0.007 $61.33\pm 0.91$ 46.52 $\pm$ 4.13
Graphormer $0.058\pm 0.015$ 1.201 $\pm$ 0.032 OOM OOM $0.067\pm 0.046$ 0.287 $\pm$ 0.007 OOM OOM
ESM-1b 0.562 $\pm$ 0.003 0.548 $\pm$ 0.004 0.682 $\pm$ 0.077 0.282 $\pm$ 0.010 0.664 $\pm$ 0.002 0.171 $\pm$ 0.001 Length Exceeded Length Exceeded
ESM-2-650M 0.548 $\pm$ 0.004 0.563 $\pm$ 0.002 0.689 $\pm$ 0.020 0.273 $\pm$ 0.014 0.694 $\pm$ 0.001 0.154 $\pm$ 0.001 71.53 $\pm$ 0.58 63.33 $\pm$ 3.68
Prot-T5-XL 0.553 $\pm$ 0.003 0.535 $\pm$ 0.006 $\bf 0.767\pm 0.008$ 0.272 $\pm$ 0.030 0.756 $\pm$ 0.003 0.136 $\pm$ 0.003 74.95 $\pm$ 1.01 62.75 $\pm$ 0.76
ProtBert 0.566 $\pm$ 0.001 0.583 $\pm$ 0.007 0.646 $\pm$ 0.064 0.322 $\pm$ 0.026 0.572 $\pm$ 0.001 0.207 $\pm$ 0.001 70.70 $\pm$ 0.47 62.53 $\pm$ 2.67
DeepProt-T5 0.614 0.642 0.725 0.551 0.874** 0.084** 77.00 68.25**
ChemLLM-7B -0.019 19.878 -0.187 46273.897 0.017 111.76 50.65 1.64
LlaSMol-Mistral-7B NaN 2.083 0.094 1.359 NaN 0.268 50.30 6.89
@@
Model Subcellular Binary
Acc $\uparrow$ Averaged F1 $\uparrow$ PR-AUC $\uparrow$ Averaged F1 $\uparrow$
@@
CNN 50.18 $\pm$ 1.21 30.44 $\pm$ 0.28 $90.88\pm 0.31$ 87.74 $\pm$ 0.08
CNN-RNN 52.58 $\pm$ 0.11 38.21 $\pm$ 0.57 91.43 $\pm$ 0.45 88.56 $\pm$ 1.03
Transformer $42.63\pm 0.68$ 25.46 $\pm$ 0.01 $78.38\pm 0.25$ 72.20 $\pm$ 2.21
GCN $47.45\pm 0.47$ 34.88 $\pm$ 0.79 $83.43\pm 0.10$ 81.80 $\pm$ 0.07
GAT $45.14\pm 0.10$ 27.03 $\pm$ 0.47 $82.15\pm 0.41$ 81.89 $\pm$ 0.19
NeuralFP $45.20\pm 0.49$ 27.07 $\pm$ 0.93 $81.14\pm 0.06$ 79.95 $\pm$ 0.09
AttentiveFP $42.38\pm 1.25$ 23.50 $\pm$ 1.34 $80.58\pm 0.30$ 80.08 $\pm$ 0.39
ESM-2-650M 79.07 $\pm$ 0.05 $6$$\pm$ 1.14 96.63 $\pm$ 0.18 91.76 $\pm$ 0.47
Prot-T5-XL 80.67 $\pm$ 0.04 69.86 $\pm$ 0.39 97.03 $\pm$ 0.13 93.48 $\pm$ 0.07
LlaSMol-Mistral-7B 15.65 4.99 57.52 0.00
ChemLLM-7B 6.24 0.59 56.32 1.43
DeepProt-T5 82.69 82.52 92.17 92.18
@@
Model PPI Affinity Yeast PPI Human PPI
$R^{2}$uparrow MAE $\downarrow$ PR-AUC $\uparrow$ Averaged F1 $\uparrow$ PR-AUC $\uparrow$ Averaged F1 $\uparrow$
@@
CNN $\bf 0.646\pm 0.003$ 1.764 $\pm$ 0.051 $51.93\pm 0.92$ 25.90 $\pm$ 0.10 $70.37\pm 1.22$ 69.65 $\pm$ 1.63
CNN-RNN 0.584 $\pm$ 0.026 1.886 $\pm$ 0.108 $53.28\pm 0.85$ 47.20 $\pm$ 3.62 $70.45\pm 2.68$ 69.87 $\pm$ 2.59
Transformer $0.425\pm 0.021$ 2.081 $\pm$ 0.133 $53.79\pm 1.07$ 51.93 $\pm$ 0.40 $59.36\pm 4.00$ 68.64 $\pm$ 1.06
GCN $0.366\pm 0.034$ 2.443 $\pm$ 0.036 58.98 $\pm$ 0.72 48.13 $\pm$ 3.88 82.21 $\pm$ 1.13 71.05 $\pm$ 2.90
GAT $0.230\pm 0.001$ 2.463 $\pm$ 0.015 $53.72\pm 0.39$ 57.00 $\pm$ 3.83 $77.63\pm 3.13$ 73.92 $\pm$ 3.50
NeuralFP $0.100\pm 0.054$ 2.555 $\pm$ 0.040 $57.00\pm 1.51$ 58.94 $\pm$ 4.74 $80.11\pm 1.25$ 67.62 $\pm$ 1.03
ESM-2-650M 0.592 $\pm$ 0.001 1.893 $\pm$ 0.005 67.36 $\pm$ 0.80 63.99 $\pm$ 1.00 96.17 $\pm$ 0.18 87.63 $\pm$ 0.60
Prot-T5-XL 0.573 $\pm$ 0.011 1.979 $\pm$ 0.007 69.84 $\pm$ 0.46 66.97 $\pm$ 0.01 95.36 $\pm$ 0.07 87.64 $\pm$ 0.98
LlaSMol-Mistral-7B -0.008 20.335 53.64 51.39 49.48 49.15
ChemLLM-7B 0.082 28.771 53.65 15.45 47.15 15.49
DeepProt-T5 0.643 1.870 63.20 66.70 76.37 80.28
Table 5: Results of epitope and paratope prediction (residue-level classification). Structure-based and pretrained protein language models took large GPU or CPU memory so we disregrad them in residue-level prediction. The same strategy is applied to secondary structure as well.
Model IEDB PDB-Jespersen SAbDab-Liberis
ROC-AUC $\uparrow$ Averaged F1 $\uparrow$ ROC-AUC $\uparrow$ Averaged F1 $\uparrow$ ROC-AUC $\uparrow$ Averaged F1 $\uparrow$
@@
Model TAP SAbDab-Chen CRISPR-Leenay
$R^{2}$uparrow MAE $\downarrow$ $R^{2}$uparrow MAE $\downarrow$ $R^{2}$uparrow MAE $\downarrow$
@@
CNN 0.469 $\pm$ 0.066 $3.217\pm 0.026$ 0.547 $\pm$ 0.014 0.219 $\pm$ 0.006 0.781 $\pm$ 0.009 0.0745 $\pm$ 0.0005
CNN-RNN 0.972 $\pm$ 0.008** 0.712 $\pm$ 0.069** 0.486 $\pm$ 0.012 $0.226\pm 0.001$ 0.771 $\pm$ 0.004 $0.0755\pm 0.0010$
Transformer 0.030 $\pm$ 0.011 $3.476\pm 0.004$ 0.452 $\pm$ 0.011 $0.238\pm 0.012$ 0.200 $\pm$ 0.054 $0.1216\pm 0.0020$
GCN 0.614 $\pm$ 0.054 $2.761\pm 0.155$ 0.434 $\pm$ 0.030 $0.326\pm 0.015$ 0.066 $\pm$ 0.010 $0.1274\pm 0.0019$
GAT 0.777 $\pm$ 0.007 $2.675\pm 0.022$ 0.356 $\pm$ 0.015 $0.310\pm 0.010$ 0.054 $\pm$ 0.010 $0.1232\pm 0.0001$
NeuralFP 0.205 $\pm$ 0.035 $3.436\pm 0.015$ 0.452 $\pm$ 0.027 $0.253\pm 0.011$ 0.177 $\pm$ 0.029 $0.1243\pm 0.0001$
ESM-2-650M 0.866 $\pm$ 0.007 2.452 $\pm$ 0.027 0.600 $\pm$ 0.001 0.224 $\pm$ 0.001 0.110 $\pm$ 0.011 0.1218 $\pm$ 0.0010
Prot-T5-XL 0.837 $\pm$ 0.010 2.417 $\pm$ 0.056 0.596 $\pm$ 0.011 0.229 $\pm$ 0.003 0.236 $\pm$ 0.004 0.1186 $\pm$ 0.0006
LlaSMol-Mistral-7B NaN 48.104 NaN 0.241 NaN 0.7922
ChemLLM-7B 0.099 45.957 -0.006 32.33 0.061 20.0503
DeepProt-T5 0.758 2.922 0.536 0.186 0.753 0.078
Table 7: Results of protein folding prediction (protein-level and residue-level classification).
Model Fold Secondary Structure
Acc $\uparrow$ Averaged F1 $\uparrow$ Acc $\uparrow$ Averaged F1 $\uparrow$
# train/valid/test 12312 / 736 / 718 8678 / 2170 / 513
CNN 8.01 $\pm$ 0.44 0.47 $\pm$ 0.11 99.44 $\pm$ 0.00 87.76 $\pm$ 0.06
CNN-RNN 7.50 $\pm$ 0.01 1.06 $\pm$ 0.01 99.93 $\pm$ 0.00 98.80 $\pm$ 0.00
Transformer 5.34 $\pm$ 0.24 0.20 $\pm$ 0.02 98.64 $\pm$ 0.00 49.66 $\pm$ 0.00
GCN 8.09 $\pm$ 0.36 2.34 $\pm$ 0.15 / /
ESM-2-650M 49.80 $\pm$ 0.36 25.55 $\pm$ 0.61 / /
Prot-T5-XL 50.28 $\pm$ 0.52 24.24 $\pm$ 0.48 / /
LlaSMol-Mistral-7B 0.00 0.00 / /
ChemLLM-7B 0.64 0.04 / /
DeepProt-T5 74.00** 73.61** / /
@@ 3.4 Results & Analysis
For each method, we used three different random seeds to conduct independent runs and reported the average results and their standard deviations. The results of protein function prediction are reported in Table 2 and Table 3. The results of protein-protein interaction are reported in Table 4. The results of epitope and paratope interaction are reported in Table 5. The results of antibody developability prediction are reported in Table 6. The results of protein structure prediction are reported in Table 7.
@@
Figure 4: We recorded training time and GPU memory assumptions for each task and each model. Specifically, we extract three representative methods: CNN, GCN and Prot-T5 from sequence-based, structure-based and pre-trained protein language models. We observed that Prot-T5 and GCN took up more GPU memory than CNN. Prot-T5, where upstream embeddings are fixed, is more efficient in training downstream tasks. Training a GCN model took more time than training a CNN or a Prot-T5 model.
Key Observations. We summarize the following key observations as takeaways.
  • pre-trained protein language models and our DeepProt-T5 are powerful compared with sequence-based and structure-based neural architectures. Sequence-based neural architectures, such as CNN, RNN, and transformer, obtain also superior performance in most protein sequence learning tasks. Specifically, in 12 out of all the 17 tasks across various protein sequence learning tasks, both sequenced-based models (CNN, RNN, Transformer) and the pre-trained protein language models (Prot-T5-XL, ESM-2-650M and DeepProt-T5) takes the top-2 position.
  • Among all the 13 GNN-solvable tasks (except residue-level classification), graph neural networks (GNN) obtain the inferior performance compared with sequenced-based and protein language models. The potential reason would be that SMILES or original string didnt provide the 3d information (coordinates) about a protein, the graph topology given by edge featurizer is ill-defined in tbe deep graph library.
  • Among all the graph neural networks (GNNs) across the whole 12 GNN-solvable tasks (except residue-level classification), the earliest variant, GCN (Kipf & Welling, 2016), achieves the best performance in 9 tasks.
  • Stability. From the learning curve (Figure 5), we find that GNN’s training curve is not stable. In contrast, the sequence-based models, including CNN, RNN, and transformer, converge more stably from the learning curve. This can be observed from Figure 5. On the contrary, training is more stable, fast and accurate for GAT when it comes to TAP dataset.
  • Computational complexity. The runtime and memory costs are reported in Figure 4. We find that GNN-based models are typically computationally inefficient. The key reason behind this is that GNN utilizes molecular graph as the feature, where each atom corresponds to a node and each chemical bond corresponds to an edge. While another model, such as CNN, RNN, and transformer, uses amino acid sequences as the input feature.